OpenStudy (anonymous):

Find the derivative of sin^2 4x^3

OpenStudy (anonymous):

I'm getting the wrong answer every time. f'(x) = (2)(sin 4x^3)(cos 4x^3)(12x^2) = (24x^2)(cos 4x^3)(sin 4x^3) That's the wrong answer, though.

OpenStudy (lgbasallote):

use the chain rule derivative of sin^2(4x^3) times derivative of 4x^3 =2sin(4x^3)(12x^2) =sin(8x^3)(12x^2)

OpenStudy (anonymous):

so you don't take the derivative of sin 4x^3?

OpenStudy (anonymous):

chain rule three times

OpenStudy (lgbasallote):

you take the derivative of sin^2 then copy 4x^3 then take the derivative of 4x^3

OpenStudy (anonymous):

Oh my God! Thanks lgbasallote! I'm studying for the test and this stuff is a couple weeks old, so I'm beginning to forget...wow!

OpenStudy (anonymous):

\[\frac{d}{dx}\sin^2(4x^3)\] not quite right still

OpenStudy (anonymous):

you have to take the derivative of a) something squared b) sin of something c) 4 something cubed

OpenStudy (anonymous):

That makes a lot of sense! Geez...I'm in trouble *gulp*...hehe.

OpenStudy (lgbasallote):

@deniatus try this site it will help you review: www.wolframalpha.com

OpenStudy (anonymous):

first the square the derivative of something squared is twice something the derivative of sine something is cosine something and the derivative of 4 something cubed is 14 something squared

OpenStudy (anonymous):

your answer should be \[2\times\sin(4x^3)\times \cos(4x^3)\times 12x^2\]

OpenStudy (anonymous):

No, it shouldn't satellite73. That's the answer I got as well and it was incorrect.

OpenStudy (anonymous):

lgbasallote is correct with his/her answer.

OpenStudy (anonymous):

well actually it is correct for the problem you asked

OpenStudy (anonymous):

Trust me, I thought so too but it isn't. I was scratching my head as to what I was doing wrong...

OpenStudy (anonymous):

it can be cleaned up as \[24x^2\sin(4x^3)\cos(4x^3)\]

OpenStudy (anonymous):

I KNOW...it's incorrect. It's the SAME answer I got each time....

OpenStudy (anonymous):

The answer is 12x^2 sin 8x^3

OpenStudy (anonymous):

you can take it to the bank, if the question is "find the derivative of \[\sin^2(4x^3)\] then the answer is \[24x^2\sin(4x^3)\cos(4x^3)\]

OpenStudy (anonymous):

well sure, that is the same thing right?

OpenStudy (anonymous):

Yup, that's exactly the problem. However, the answer is different in the back of the book.

OpenStudy (anonymous):

to get from one to the other is the double angle formula

OpenStudy (anonymous):

they are using the fact that \[\sin(2x)=2\sin(x)\cos(x)\]

OpenStudy (anonymous):

it is the same thing exactly, just used a trig identity to rewrite it and confuse you

OpenStudy (anonymous):

in other words \[2\sin(4x^3)\cos(4x^3)=\sin(8x^3)\] by the double angle formula

OpenStudy (anonymous):

Ah so you and I are correct then...

OpenStudy (anonymous):

yes of course they are both the same, but trust me you will not get to the answer in the back of hte book directly. you have to first write it out and then use double angle formula